OpenStudy (anonymous):

Given f(x) = 3x + 2 and g(x) = x^2 + 5x - 2, find f(x) - g(x).

OpenStudy (welshfella):

3x + 2 - (x^2 + 5x - 2)

OpenStudy (anonymous):

can u help

OpenStudy (marihelenh):

0-x^2=-x^2 3x-5x=-2x 2-(-2)=2+2=4 Your final answer is x^2-2x+4.